OpenStudy (anonymous):

If y = 12x - 2 were changed to y = 12x, how would the graph of the new function compare with the original?

OpenStudy (anonymous):

you would shifted 2 units down, (or if you want y = 12x - 2 → y=2(6x-1) y = 12x y=2(6x) thus one unit to the right.

OpenStudy (anonymous):

I mean two units up, because y=12x-2 is the initial function.

OpenStudy (anonymous):

OR one unit to the left (not one unit to the right)

OpenStudy (anonymous):

thank you i really needed your help

OpenStudy (anonymous):

anytime .....